OpenStudy (anonymous):

what are the x- intercepts of the graph of y + 20 = x^2 - x

OpenStudy (ivettef365):

when y=0 you have 0 + 20 = x^2 - x so you have 0 = x^2 -x -20 can you solve this

OpenStudy (anonymous):

Does It factor to (x-5) (x+4) so the solutions are 5 and -4?

OpenStudy (ivettef365):

that is right when x = 5 and when x = -4

OpenStudy (anonymous):

Thank you

OpenStudy (ivettef365):

yw :)